An object in free fall from

50 m high, so the height is

when t is h(t) = 50 – 4.9t2

.

What is the speed at

t = 2?

2. Determine the equation of the tangent line

on the curve y = x3 at the point (2.8)


1
Expert's answer
2021-11-03T19:05:34-0400

"c\displaystyle
\text{The speed is given by the derivative of h(t)}\\
h(t) = -4.9t^2 + 50\\
h'(t) = -9.4t\\
\text{At t = 2,$ h'(t)$ = -18.8}\\
\text{For the 2nd question, the equation of the tangent is given by $y= mx+c$}\\
\text{where m represents the gradient and c the intercept, we find m by computing the }\\
\text{the derivative}\\
y'(x) =3x^2, \text{at x = 2.8, $y = 3(2.8)^2 =23.52$}\\
\text{Next, we compute the intercept c, at x = 2.8}\\
2.8^3 = 23.52(2.8) + c\\
\implies c = -43.904\\
\text{Thus, the equation of the tangent is given by}\\
y = 23.52x -43.904"